# Stochastic Polyak Step-sizes and Momentum: Convergence Guarantees and Practical Performance

### v1

@@ 3 Convergence Analysis
In this section, we present the convergence results for SHB with all of the proposed Polyak step-sizes. For the formal definitions and helpful lemmas, see Appendix C. The proofs can be found in Appendix D. For all of our results, we make the assumption $\sigma^{2}:=\operatorname{\mathbb{E}}_{S_{t}}[f_{S_{t}}(x^{*})-\ell_{S_{t}}^{*}]=f(x^{*})-\operatorname{\mathbb{E}}_{S_{t}}[\ell_{S_{t}}^{*}]<\infty$. This assumption was first introduced in Loizou et al. (2021) and Orvieto et al. (2022). Note that $\sigma^{2}<\infty$ when all $f_{i}$ are lower bounded like we have assumed. We say that problem (1) is interpolated if $\sigma^{2}=0$. If the interpolation condition is satisfied then there exists $x^{*}\in X^{*}$ such that $f(x^{*})=f_{S_{t}}(x^{*})=\ell_{S_{t}}^{*}$ for all $S_{t}\subseteq[n]$. Many modern machine learning models satisfy this condition. Examples include non-parametric regression Liang and Rakhlin (2020) and over-parameterized deep neural networks (Zhang et al., 2021; Ma et al., 2018).
@@ 3.1 Convergence to a neighborhood of the solution
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of SHB with MomSPSmax with $c=1$ and $\beta\in\left[0,\frac{\alpha}{2\gamma_{b}-\alpha}\right)$ where $\alpha=\min\left\{\frac{1}{2L_{\max}},\gamma_{b}\right\}$ and $L_{\max}=\max_{i\in[n]}\{L_{i}\}$, converge as
@@
In addition, note that in Theorem 3.1, there is a restriction in the momentum coefficient. This stems from the technique used in the proof (which forces $C_{1}$ and $C_{2}$ to be positive). However, this restriction is not vital as in practical scenarios, the method converges even when $\beta$ is selected outside the given interval. Moreover, let us highlight that the constants $C_{1}$ and $C_{2}$ are increasing when viewed as functions of $\beta$ in the given interval (see also Figure 14). This means that our theory suggests that $\beta=0$ (no momentum) is the best theoretical choice, which is typical in many works on stochastic methods with momentum (Wang et al., 2023; Loizou and Richtárik, 2020). This is not ideal but does not undermine the importance of Theorem 3.1, as this is the first result showing the convergence of SHB with SPS in a non-interpolated setting.
@@
Corollary 3.2.
@@
Note that in the setting of Corollary 3.2, we have used $f(x^{*})$ instead of a lower bound $\ell^{*}$. This is in accordance with previous work in the deterministic setting, (Polyak, 1987). We highlight that in Hazan and Kakade (2019) (PS) and Wang et al. (2023) (ALR-MAG), there is no upper bound in the step-size for the deterministic setting. This is due to the fact that in these works, the step-size is derived as a minimizer of an upper bound of the quantity $\|x^{t}-x^{*}\|^{2}$. Our step-size MomPSmax is not a minimizer of such bound when $\beta>0$. Nevertheless, when $\beta=0$ (no momentum), by setting $\gamma_{b}=\infty$ we recover the corresponding result in Hazan and Kakade (2019). In our setting, we cannot set $\gamma_{b}=\infty$ when $\beta>0$ because the term $\beta\gamma_{b}$ appears in the denominator of our convergence result.
Corollary 3.3.
@@
Assume interpolation ($\sigma^{2}=0$) and let all assumptions of Theorem 3.1 be satisfied. Then SHB with MomSPSmax and $\beta\in[0,\alpha/(2\gamma_{b}-\alpha))$ where $\alpha=\min\left\{\frac{1}{2L_{\max}},\gamma_{b}\right\}$ converges as
$\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]\leq\frac{(1-\beta)\|x^{0}-x^{*}\|^{2}}{(\alpha\beta+\alpha-2\beta\gamma_{b})T},$
where $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T-1}x^{t}$.
@@
In the closely related works Loizou et al. (2021) (SPSmax) and Wang et al. (2023) (ALR-SMAG), there is no upper bound in the proposed step-sizes for the interpolated setting. This is because these works treat the interpolated regime the same as the deterministic regime (the proofs of the results in these two settings are almost identical). In our analysis, we focus on the fully stochastic setting (interpolation is only a corollary), and the result of Corollary 3.3 shows a saddle connection between the momentum parameter and the parameter $\gamma_{b}$ (we cannot select $\gamma_{b}=\infty$ when momentum is used). Nevertheless, via our corollary, when $\beta=0$ by setting $\gamma_{b}=\infty$, we can recover the corresponding result in Loizou et al. (2021).
@@
Another interesting outcome of Theorem 3.1 is a novel analysis for SHB with a constant step size.
Corollary 3.4.
@@
Let all assumptions of Theorem 3.1 be satisfied. If $\gamma_{b}\leq\frac{1}{2L_{\max}}$, then SHB with $\gamma\leq\frac{1-\beta}{2L_{\max}}$ and $\beta\in[0,1)$ converges as
$\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]\leq\frac{\|x^{0}-x^{*}\|^{2}}{T\gamma_{b}}+\sigma^{2},$
where $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T-1}x^{t}$.
@@
Note that in Corollary 3.4, there is no restriction that depends on the smoothness parameter in the momentum parameter $\beta$. We allow to have $\beta\in[0,1)$. A similar result is obtained in Liu et al. (2020), where the authors establish the convergence rate $O\left(\frac{f(x^{0})-f(x^{*})}{T}+\frac{L\gamma\hat{\sigma}^{2}}{1-\beta}\right)$ for the quantity $\frac{1}{T}\sum_{t=0}^{T-1}\operatorname{\mathbb{E}}[\|\nabla f_{S_{t}}(x^{t})\|^{2}]$ when $\gamma\leq\frac{(1-\beta)^{2}}{L}\min\left\{\frac{1}{4-\beta+\beta^{2}},\frac{1}{2\sqrt{2\beta+2\beta^{2}}}\right\}$. Here $\hat{\sigma}^{2}$ is assumed to be a bound of the variance $\operatorname{\mathbb{E}}_{S_{t}}\|\nabla f_{S_{t}}(x^{t})-\nabla f(x^{t})\|^{2}\leq\hat{\sigma}^{2}$. In comparison, in our analysis, we provide the same asymptotic rate $O(1/T)$ for the common in the convex setting quantity $\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]$. By comparing the two results, our step size is larger when $\beta\geq\sqrt{5}-2\approx 0.236$. However, in our result, the neighborhood of convergence is constant and does not depend on the step-size $\gamma$ or the momentum coefficient. For a numerical comparison of SHB with the two different constant step-sizes, see Section E.2.
@@ 3.2 Convergence to the exact solution
In this section, we provide theoretical results for the adaptive decreasing step-sizes, MomDecSPS and MomAdaSPS. The main advantage of these decreasing step-sizes is that we can guarantee convergence to the exact solution while keeping the main adaptiveness properties. Due to the nature of the decreasing step sizes, there is no restriction in the momentum parameter, as was the case with Theorem 3.1. For the results of this section, we make the extra assumption of bounded iterates, i.e., we assume that $D^{2}=\max_{t\in[T]}\|x^{t}-x^{*}\|^{2}$ is finite. This is a standard assumption for several adaptive step-sizes, see: (Reddi et al., 2018; Ward et al., 2020; Orvieto et al., 2022; Jiang and Stich, 2023).
@@
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of SHB with MomDecSPS with $c=1$ and $\beta\in[0,1)$ converge as:
$\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]\leq\frac{2\beta[f(x^{0})-f(x^{*})]}{(1-\beta)T}+\frac{(1+\beta)D^{2}}{(1-\beta)\alpha\sqrt{T}}+\frac{2\sigma^{2}}{\sqrt{T}},$
where $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T-1}x^{t}$, $\alpha=\min\left\{\frac{1}{2L},\gamma_{b}\right\}$ and $D^{2}=\max_{t\in[T]}\|x^{t}-x^{*}\|^{2}$.
@@ 4 Numerical Experiments
4.1 Deterministic Setting
For the deterministic setting, we focus on the least squares problem. The loss function in this case is given by $f(x)=\frac{1}{2}\|Ax-b\|^{2}$, where $A\in\mathbb{R}^{n\times d}$ and $b\in\mathbb{R}^{n}$. In our experiments, we follow the setting of Wang et al. (2023) and choose $n=d=1000$ while the matrix $A$ has been generated according to Lenard and Minkoff (1984) such that the condition number of $A^{T}A$ is $10^{4}$. We test the following algorithms: Gradient Descent (GD), Nesterov’s Accelerated Gradient Descent (AGD) (Nesterov, 2018), Gradient Descent with Polyak’s step-size (Polyak, 1987), Adaptive Heavy Ball (AHB) (Saab Jr et al., 2022), Polyak’s Heavy Ball (HB) (Polyak, 1964), Accelerated Gradient Method (AGM) (Barré et al., 2020), ALR-MAG (Wang et al., 2023) and finally our step-size from Corollary 3.2, called (MomPSmax), which guarantees $O(1/T)$ rate. For the least squares problem, there is an established theory from Polyak (1964) that guarantees acceleration of HB with optimal step-size and momentum coefficient given by $\beta^{*}=(\sqrt{L}-\sqrt{\mu})^{2}/(\sqrt{L}+\sqrt{\mu})^{2}$ and $\gamma^{*}=(1+\sqrt{\beta^{*}})/L$. For ALR-MAG and SHB with MomPSmax, we have used the optimal momentum $\beta^{*}$ for direct comparison with optimal HB. Moreover, for GD Polyak, ALR-MAG, and MomPSmax, we have used the precomputed $f(x^{*})$ (via GD), and for MomPSmax, we have chosen $\gamma_{b}=100$ (for other choices, see Section E.4). The results in Figure 2 show that MomPSmax has performance comparable to AGM and ALR-MAG which are faster than vanilla Polyak step-size after the initial iterations. Furthermore, to explore the behavior of the various adaptive step-sizes, we plot the Polyak, the ALR-SMAG, the MomPSmax, and the optimal HB step-sizes. We can see that the original Polyak step-size oscillates around the optimal HB step-size while MomPSmax converges from below to the optimal.
Figure 2: Comparison of MomPSmax against various deterministic algorithms for consistent least squares. Left: Relative error, Right: Step-sizes.
4.2 Stochastic Setting: Convex Problems
Here, we compare our step-sizes with previous works in the stochastic setting. As noticed initially for SPSmax in Loizou et al. (2021), the value of the upper bound $\gamma_{b}$ that results in good convergence depends on the problem and requires careful parameter tuning. To alleviate this problem, Loizou et al. (2021) uses a smoothing procedure that prevents large fluctuations in the step-size across iterations. That is, $\gamma_{b}^{t}=\tau^{b/n}\gamma_{t-1}$ for each iteration $t$ where $\tau=2$, $b$ is the batch-size, and $n$ is the number of examples. We use the same smoothing trick for $\gamma_{b}$ in the implementation of the proposed methods. Similar smoothing procedures have been used in (Tan et al., 2016; Vaswani et al., 2019b).
@@ 4.2 Stochastic Setting: Convex Problems
Figure 3: Comparison between our proposed MomSPSmax, other momentum methods, SPSmax (no momentum) and constant step-size SGD. Setting: Unregularized logistic regression, LIBSVM dataset: vowel, Batch-size for all methods: 52
Figure 4: Comparison of our proposed MomDecSPS and MomAdaSPS between DecSPS and AdaSPS as well as AdaGrad-Norm. Setting: Unregularized logistic regression, LIBSVM Dataset: letter, Batch-size for all methods: 1500
4.3 MomSPSmax for DNNs
In our final experiment, we go beyond the convex setting of our convergence guarantees, and we test MomSPSmax on the training of DNNs. We consider non-convex minimization for multi-class classification using deep network models on the CIFAR-10 and CIFAR-100 datasets, (Krizhevsky et al., 2009). We use two standard image-classification architectures: ResNet34 (He et al., 2016) and DenseNet121 (Huang et al., 2017). For space concerns, we report only the ResNet experiments in the main paper and relegate the DenseNet to the Appendix. We present the results for CIFAR-10 in Figure 5 and for CIFAR-100 in Figure 6. We observe that MomSPSmax consistently outperforms its no-momentum counterpart SPSmax and has competitive generalization performance compared to other popular optimizers. An important observation is that for these experiments (CIFAR-10 and CIFAR-100 with ResNet34), the closely related methods that combine stochastic Polyak step-size with momentum (ALR-SMAG from Wang et al. (2023) and SPSmax with naive momentum from Section 2.2) diverge and not generalize well. This further highlights the importance of balancing the stochastic Polyak step size and the momentum parameter we provide in Section 2.3.
Figure 5: Comparison between our proposed MomSPSmax, other momentum methods, SPSmax (no momentum) and constant step-size SGD. Model: ResNet 34, Dataset: CIFAR-10, Batch-size for all methods: 256
Figure 6: Comparison between our proposed MomSPSmax, other momentum methods, SPSmax (no momentum) and constant step-size SGD. Model: ResNet 34, Dataset: CIFAR-100, Batch-size for all methods: 256
5 Conclusion and Future Work
We proposed and theoretically analyzed three novel adaptive Polyak step-size selections for SHB: MomSPSmax, MomDecSPS and MomAdaSPS. We provided convergence guarantees for the three variants for solving convex and smooth optimization problems and verified our theory via numerical experiments. There are many possible interesting extensions of our work: explore the convergence of the proposed methods for non-convex or non-smooth problems, work on adaptive Polyak-type momentum algorithms that can guarantee faster convergence over the no-momentum counterparts, and study extensions in distributed and decentralized settings.
@@ 5 Conclusion and Future Work
The Supplementary Material is organized as follows: In Appendix A, we have more details on related work on adaptive and momentum methods. In Appendix B, we include the main pseudo-codes for our algorithms. In Appendix C, we give the basic definitions and lemmas as well as the basic theory of IMA. Appendix D presents the proofs of the theoretical guarantees from the main paper. In Appendix E, we describe in detail our experimental setup and provide additional experiments.

### v2

@@ 3 Convergence Analysis
In this section, we present the convergence results for SHB with all of the proposed Polyak step-sizes. For the formal definitions and helpful lemmas, see Appendix C. The proofs can be found in Appendix D. For all of our results, we make the following assumption:
Assumption 3.1 (Finite optimal objective difference).
$\sigma^{2}:=\operatorname{\mathbb{E}}_{S_{t}}[f_{S_{t}}(x^{*})-\ell_{S_{t}}^{*}]=f(x^{*})-\operatorname{\mathbb{E}}_{S_{t}}[\ell_{S_{t}}^{*}]<\infty$
This assumption was first introduced in Loizou et al. (2021) and Orvieto et al. (2022). Note that $\sigma^{2}<\infty$ when all $f_{i}$ are lower bounded like we have assumed. We say that problem (1) is interpolated if $\sigma^{2}=0$. If the interpolation condition is satisfied then there exists $x^{*}\in X^{*}$ such that $f(x^{*})=f_{S_{t}}(x^{*})=\ell_{S_{t}}^{*}$ for all $S_{t}\subseteq[n]$. Many modern machine learning models satisfy this condition. Examples include non-parametric regression (Liang and Rakhlin, 2020) and over-parameterized deep neural networks (Zhang et al., 2021; Ma et al., 2018).
@@ 3.1 Convergence to a neighborhood of the solution
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of SHB with MomSPSmax with $c=1$ and $\beta\in\left[0,\frac{\alpha}{2\gamma_{b}-\alpha}\right)$ where $\alpha=\min\left\{\frac{1}{2L_{\max}},\gamma_{b}\right\}$ and $L_{\max}=\max_{i}\{L_{i}\}$, converge as
@@
In addition, note that in Theorem 3.2, there is a restriction in the momentum coefficient. This stems from the technique used in the proof (which forces $C_{1}$ and $C_{2}$ to be positive). However, this restriction is not vital as in practical scenarios, the method converges even when $\beta$ is selected outside the given interval.
Moreover, let us highlight that the constants $C_{1}$ and $C_{2}$ are increasing when viewed as functions of $\beta$ in the given interval (see also Figure 14). This means that our theory suggests that $\beta=0$ (no momentum) is the best theoretical choice, which is typical in many works on stochastic methods with momentum (Wang et al., 2023; Loizou and Richtárik, 2020). This is not ideal but does not undermine the importance of Theorem 3.2, as this is the first result showing the convergence of SHB with SPS in a non-interpolated setting.
@@
Corollary 3.3 (Deterministic Heavy Ball Method).
@@
Note that in the setting of Corollary 3.3, we have used $f(x^{*})$ instead of a lower bound $\ell^{*}$. This is in accordance with previous work in the deterministic setting (Polyak, 1987). We highlight that in Hazan and Kakade (2019) (PS) and Wang et al. (2023) (ALR-MAG), there is no upper bound in the step-size for the deterministic setting. This is due to the fact that in these works, the step-size is derived as a minimizer of an upper bound of the quantity $\|x^{t}-x^{*}\|^{2}$. Our step-size MomPSmax is not a minimizer of such bound when $\beta>0$. Nevertheless, when $\beta=0$ (no momentum), by setting $\gamma_{b}=\infty$ we recover the corresponding result in Hazan and Kakade (2019). In our setting, we cannot set $\gamma_{b}=\infty$ when $\beta>0$ because the term $\beta\gamma_{b}$ appears in the denominator of our convergence result.
Corollary 3.4 (Interpolation).
@@
Assume interpolation ($\sigma^{2}=0$) and let all assumptions of Theorem 3.2 be satisfied. Then SHB with MomSPSmax and $\beta\in[0,\alpha/(2\gamma_{b}-\alpha))$ where $\alpha=\min\left\{\frac{1}{2L_{\max}},\gamma_{b}\right\}$ converges as $\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]\leq\frac{(1-\beta)\|x^{0}-x^{*}\|^{2}}{(\alpha\beta+\alpha-2\beta\gamma_{b})T}$, where $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T-1}x^{t}$.
@@
In the closely related works Loizou et al. (2021) (SPSmax) and Wang et al. (2023) (ALR-SMAG), there is no upper bound in the proposed step-sizes for the interpolated setting. In our analysis, we focus on the fully stochastic setting (interpolation is only a corollary), and the result of Corollary 3.4 shows a saddle connection between the momentum parameter and the parameter $\gamma_{b}$ (we cannot select $\gamma_{b}=\infty$ when momentum is used). Nevertheless, via our corollary, when $\beta=0$ by setting $\gamma_{b}=\infty$, we can recover the corresponding result in Loizou et al. (2021).
@@
Another Corollary of Theorem 3.2 is a novel analysis for SHB with constant step-size.
Corollary 3.5 (Constant Step-size).
@@
Let all assumptions of Theorem 3.2 be satisfied. If $\gamma_{b}\leq\frac{1}{2L_{\max}}$, then SHB with $\gamma\leq\frac{1-\beta}{2L_{\max}}$ and $\beta\in[0,1)$ converges as $\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]\leq\frac{\|x^{0}-x^{*}\|^{2}}{T\gamma}+\sigma^{2}$, where $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T-1}x^{t}$.
@@
Note that in Corollary 3.5, there is no restriction that depends on the smoothness parameter in the momentum parameter $\beta$. We allow to have $\beta\in[0,1)$. A similar result is obtained in Liu et al. (2020), where the authors establish the convergence rate $O\left(\frac{f(x^{0})-f(x^{*})}{T}+\frac{L\gamma\hat{\sigma}^{2}}{1-\beta}\right)$ for the quantity $\frac{1}{T}\sum_{t=0}^{T-1}\operatorname{\mathbb{E}}[\|\nabla f_{S_{t}}(x^{t})\|^{2}]$ when $\gamma\leq\frac{(1-\beta)^{2}}{L}\min\left\{\frac{1}{4-\beta+\beta^{2}},\frac{1}{2\sqrt{2\beta+2\beta^{2}}}\right\}$. Here $\hat{\sigma}^{2}$ is assumed to be a bound of the variance $\operatorname{\mathbb{E}}_{t}\|\nabla f_{S_{t}}(x^{t})-\nabla f(x^{t})\|^{2}\leq\hat{\sigma}^{2}$. In comparison, in our analysis, we provide the same asymptotic rate $O(1/T)$ for the common in the convex setting quantity $\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]$. By comparing the two results, our step-size is larger when $\beta\geq\sqrt{5}-2\approx 0.236$. However, in our result, the neighborhood of convergence is constant and does not depend on the step-size $\gamma$ or the momentum coefficient. For a numerical comparison of SHB with the two different constant step-sizes, see Section F.2.
@@ 3.2 Convergence to the exact solution
In this section, we provide theoretical results for the adaptive decreasing step-sizes, MomDecSPS and MomAdaSPS. The main advantage of these decreasing step-sizes is that we can guarantee convergence to the exact solution while keeping the main adaptiveness properties. Due to the nature of the decreasing step-sizes, there is no restriction in the momentum parameter, as was the case with Theorem 3.2. For the results of this section, we make the extra assumption of bounded iterates, i.e., we assume that $D^{2}=\max_{t\in[T]}\|x^{t}-x^{*}\|^{2}$ is finite. This is a standard assumption for several adaptive step-sizes, see: (Reddi et al., 2018; Ward et al., 2020; Orvieto et al., 2022; Jiang and Stich, 2023).
@@
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Let $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T-1}x^{t}$, $\alpha=\min\left\{\frac{1}{2L},\gamma_{b}\right\}$ and $D^{2}=\max_{t\in[T]}\|x^{t}-x^{*}\|^{2}$. Then, the iterates of SHB with MomDecSPS with $c=1$ and $\beta\in[0,1)$ converge as:
$\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]\leq\frac{2\beta[f(x^{0})-f(x^{*})]}{(1-\beta)T}+\frac{(1+\beta)D^{2}}{(1-\beta)\alpha\sqrt{T}}+\frac{2\sigma^{2}}{\sqrt{T}}.$
@@ 4 Numerical Experiments
Figure 2: Comparison of deterministic algorithms for least squares. Left: Relative error, Right: Step-sizes.
Deterministic Setting. For the deterministic setting, we focus on the least squares problem. The loss function in this case is given by $f(x)=\frac{1}{2}\|Ax-b\|^{2}$, where $A\in\mathbb{R}^{n\times d}$ and $b\in\mathbb{R}^{n}$. In our experiments, we follow the setting of Wang et al. (2023) and choose $n=d=1000$ while the matrix $A$ has been generated according to Lenard and Minkoff (1984) such that the condition number of $A^{T}A$ is $10^{4}$. We test the following algorithms: Gradient Descent (GD), Nesterov’s Accelerated Gradient Descent (AGD) (Nesterov, 2018), Gradient Descent with Polyak’s step-size (Polyak, 1987), Adaptive Heavy Ball (AHB) (Saab Jr et al., 2022), Polyak’s Heavy Ball (HB) (Polyak, 1964), Accelerated Gradient Method (AGM) (Barré et al., 2020), ALR-MAG (Wang et al., 2023) and finally our step-size from Corollary 3.3, called (MomPSmax), which guarantees $O(1/T)$ rate. For the least squares problem, there is an established theory from Polyak (1964) that guarantees acceleration of HB with optimal step-size and momentum coefficient given by $\beta^{*}=(\sqrt{L}-\sqrt{\mu})^{2}/(\sqrt{L}+\sqrt{\mu})^{2}$ and $\gamma^{*}=(1+\sqrt{\beta^{*}})/L$. For ALR-MAG and SHB with MomPSmax, we have used the optimal momentum $\beta^{*}$ for direct comparison with optimal HB. Moreover, for GD Polyak, ALR-MAG, and MomPSmax, we have used the precomputed $f(x^{*})$ (via GD), and for MomPSmax, we have chosen $\gamma_{b}=100$ (for other choices, see Section F.4). The results in Figure 2 show that MomPSmax has performance comparable to AGM and ALR-MAG which are faster than vanilla Polyak step-size after the initial iterations. Furthermore, to explore the behavior of the various adaptive step-sizes, we plot the Polyak, the ALR-SMAG, the MomPSmax, and the optimal HB step-sizes. We can see that the original Polyak step-size oscillates around the optimal HB step-size while MomPSmax converges from below to the optimal.
Stochastic Setting: Convex Problems. Here, we compare our step-sizes with previous works in the stochastic setting. As noticed initially for SPSmax (Loizou et al., 2021), the value of the upper bound $\gamma_{b}$ that results in good convergence depends on the problem and requires careful parameter tuning. To alleviate this problem, Loizou et al. (2021) uses a smoothing procedure that prevents large fluctuations in the step-size across iterations. That is, $\gamma_{b}^{t}=\tau^{b/n}\gamma_{t-1}$ for each iteration $t$ where $\tau=2$, $b$ is the batch-size, and $n$ is the number of examples. We use the same smoothing trick for $\gamma_{b}$ in the implementation of the proposed methods. Similar smoothing procedures have been used in Tan et al. (2016); Vaswani et al. (2019b).
@@
Figure 3: LibSVM dataset: vowel, Batch size: 52
Figure 4: LibSVM dataset: letter, Batch size: 1500
MomSPS for DNNs. In our final experiment, we go beyond the convex setting of our convergence guarantees, and we test MomSPSmax on the training of DNNs. We consider non-convex minimization for multi-class classification using deep network models on the CIFAR 10 and CIFAR 100 datasets (Krizhevsky et al., 2009). We use two standard image-classification architectures: ResNet18 and ResNet34, (He et al., 2016). For space concerns, we report only the ResNet34 experiments in the main paper and relegate the ResNet18 to the Appendix. In these experiments we run MomSPSmax with $c=0.4$. We present the results for CIFAR 10 in Figure 5 and for CIFAR 100 in Figure 6. We observe that MomSPSmax consistently outperforms its no-momentum counterpart SPSmax and has competitive generalization performance compared to other popular optimizers.
Figure 5: Resnet 34 on CIFAR 10
Figure 6: ResNet 34 on CIFAR 100
@@
The Supplementary Material is organized as follows: In Appendix A, we have more details on related work on adaptive and momentum methods. In Appendix B, we include the main pseudo-codes for our algorithms. In Appendix C, we give the basic definitions and lemmas as well as the basic theory of IMA. Appendix D presents the proofs of the theoretical guarantees from the main paper. In Appendix F, we describe in detail our experimental setup and provide additional experiments.